# CoVR: Learning Composed Video Retrieval from Web Video Captions

### v3

@@ 3.2 Our Resulting WebVid-CoVR Dataset
WebVid-CoVR: a large-scale CoVR training dataset. By applying the previously described pipeline to the WebVid2M dataset (Bain et al. 2021), we generate WebVid-CoVR, a dataset containing 1.6M CoVR triplets, which is significantly larger than prior datasets (see Table 1). On average, a video lasts 16.8 seconds, a modification text contains 4.8 words, and one target video is associated with 12.7 triplets. We study the effect of the modification text length in Section C.6 of the Appendix. WebVid-CoVR is highly diverse with 131k distinct videos and 467k distinct modification texts. Examples of CoVR triplets from the WebVid-CoVR dataset are illustrated in Figure 3. These examples (along with additional ones included in Section D.3 of the Appendix) demonstrate the diversity present in WebVid-CoVR, highlighting a wide range of content and variations in the modification texts. However, it is important to acknowledge that some noise naturally exists in the dataset, as shown in the bottom example of Figure 3, where the text does not describe the difference between the two videos due to both videos describing beautiful fields. We provide further analysis such as removal of inappropriate content, and dataset statistics of WebVid-CoVR in Section A of the Appendix.
WebVid-CoVR-Test: a new CoVR evaluation benchmark. Due to the noise in WebVid-CoVR, we manually annotate a small test set, dubbed WebVid-CoVR-Test, for evaluation. For this, we first repeat the data generation procedure described in Section 3.1, but on a different corpus of video-caption pairs. Specifically, we consider video-caption pairs from the WebVid10M corpus (Bain et al. 2021) that are not included in the WebVid2M dataset, resulting in a pool of 8 million video-caption pairs. This ensures that other models using WebVid2M for pretraining have not been exposed to any of the test examples. In the video pairs filtering stage, for each pair of captions, we here only keep one pair of videos (the one with the highest visual similarity). This results in 163k candidate triplets that could be used for testing purposes. We randomly sample 7k triplets that we use for validation and randomly sample 3.2k other triplets that we manually annotate as described below.
@@ 3.3 Training on WebVid-CoVR
CoVR-BLIP model architecture. Our model architecture builds on a pretrained image-text model, BLIP (Li et al. 2022b). BLIP is pretrained on a large dataset of image-caption pairs with three vision-language objectives: image-text contrastive learning, image-text matching, and image-conditioned language modeling. However, BLIP is not trained for composed visual retrieval with both visual and text inputs. Therefore we adapt BLIP to the CoIR/CoVR task as follows.
@@
Training. In order to train on WebVid-CoVR, we use a contrastive learning approach (van den Oord, Li, and Vinyals 2018; Radenovic et al. 2023), as it has been shown to be effective to learn strong multi-modal representations from large-scale noisy data (Radford et al. 2021). We make the following design choices. First, we create a training batch by sampling distinct target videos; and for each target video, we randomly sample an associated image-text query pair. Iterating over videos ensures that the same target video appears only once in a batch and maximizes the number of different target videos that can be used as negatives in contrastive learning. We show the benefit of this approach in Section 4.4 (Table 6). Second, we employ HN-NCE (Radenovic et al. 2023) which increases the weight of most similar samples and uses as negatives all target videos $v_{j\in\mathcal{B}}$ in the batch $\mathcal{B}$. Formally, given a training batch $\mathcal{B}$ of triplets $(q_{i},t_{i},v_{i})$, we minimize the following loss:
$\mathcal{L}(\mathcal{B})=$ $-\sum_{i\in\mathcal{B}}\text{log}\left(\frac{e^{S_{i,i}/\tau}}{\alpha\cdot{e^{S_{i,i}/\tau}}+\sum_{j\neq i}{e^{S_{i,j}/\tau}w_{i,j}}}\right)$
$-\sum_{i\in\mathcal{B}}\text{log}\left(\frac{e^{S_{i,i}/\tau}}{\alpha\cdot{e^{S_{i,i}/\tau}}+\sum_{j\neq i}{e^{S_{j,i}/\tau}w_{j,i}}}\right)$
where $\alpha$ is set to 1, temperature $\tau$ is set to $0.07$, $S_{i,j}$ is the cosine similarity between the multi-modal embedding $f_{i}$ and the target video embedding $\hat{v}_{j}$, and $w_{i,j}$ is set as in (Radenovic et al. 2023) with $\beta=0.5$.
@@ 4 Experiments
We first describe the experimental protocol including the datasets, evaluation metrics, and implementation details (Section 4.1). We then present the results of CoVR on our new video benchmark (Section 4.2), as well as transfer results of CoIR on standard image benchmarks (Section 4.3). Finally, we provide ablations on our key components (Section 4.4).
@@ 4.1 Experimental Setup
Datasets. WebVid-CoVR is our proposed training CoVR dataset, and WebVid-CoVR-Test is our new CoVR benchmark, both presented in Section 3.2.
CIRR (Liu et al. 2021) is a manually annotated CoIR dataset that contains open-domain natural images from NLVR2 (Suhr et al. 2019), comprising 36.5k queries annotated on 19k images. CIRR includes two evaluation protocols: a standard one with the entire validation set as the search gallery, and a fine-grained subset, where the search space is a subgroup of six images similar to the query image (based on pretrained ResNet15 feature distance). The dataset is divided into training, validation, and testing splits with 28225/16742, 4181/2265 and 4148/2178 queries/images, respectively.
FashionIQ (Wu et al. 2021) is another CoIR dataset that contains images of fashion products, divided into three categories of Shirts, Dresses, and Tops/Tees. The query and target images were automatically paired based on title similarities (crawled from the web), and modification texts were then manually annotated. This dataset consists of 30k queries annotated on 40.5k different images. It is divided into training and validation splits with 18000/45429 and 6016/15415 queries/images, respectively.
Evaluation metrics. Following standard evaluation protocols (Liu et al. 2021), we report the video retrieval recall at rank 1, 5, 10, and 50. Recall at rank k (R@k) quantifies the number of times the correct video is among the top k results. MeanR denotes the average of R@1, R@5, R@10, and R@50. Higher recall means better performance.
@@
modalities Fusion Backbone #frames R@1 R@5 R@10 R@50
Random - - - - 00.08 00.23 00.35 01.76
Text - BLIP - 19.68 37.09 45.85 65.14
Not finetuned on Visual - BLIP 15 34.90 59.23 68.04 85.95
WebVid-CoVR Visual + Text Avg CLIP 15 44.37 69.13 77.62 93.00
Visual + Text Avg BLIP 15 45.46 70.46 79.54 93.27
Text - BLIP - 23.67 45.89 55.13 77.03
Visual - BLIP 15 38.89 64.98 74.02 92.06
Finetuned on Visual + Text MLP CLIP 1 50.55 77.11 85.05 96.60
WebVid-CoVR Visual + Text MLP BLIP 1 50.63 74.80 83.37 95.54
Visual + Text CA BLIP 1 51.80 78.29 85.84 97.07
Visual + Text CA BLIP 15 53.13 79.93 86.85 97.69
Table 2: Benchmarking on the WebVid-CoVR-Test set: We observe that using both the visual and text input modalities performs better than individual modalities alone, both with/without finetuning on WebVid-CoVR (shown at the top/bottom of the table, respectively). When using pretraining models without finetuning, we apply average fusion (Avg) for the embeddings. BLIP performs slightly better than CLIP on this benchmark. Finetuning on WebVid-CoVR brings significant benefits. In this case, fusing with the pretrained cross-attention (CA) from BLIP is more effective than training a randomly-initialized MLP fusion as done in (Baldrati et al. 2022). Moreover, using multiple frames to embed the target video brings further improvements over using the middle frame. The first row represents the random performance.
Implementation details. For our MTG-LLM, we use LLaMA 7B model (Touvron et al. 2023) that we finetune for one epoch with an initial learning rate of $3\mathrm{e}{-5}$. For our CoVR model, we use the BLIP model with ViT-L (Dosovitskiy et al. 2021) at 384 pixels finetuned for text-image retrieval on COCO and freeze the ViT for computational efficiency. We train our CoVR model on WebVid-CoVR for 4 epochs with a batch size of 2048 and an initial learning rate of $1\mathrm{e}{-5}$. To finetune on CIRR/FashionIQ, we train for 6 epochs with a batch size of 2048/1024 and an initial learning rate of $1\mathrm{e}{-4}$. We set hyperparameters based on the validation curve of WebVid-CoVR. Experiments are conducted on 4 NVIDIA A100-SXM4-80GB GPUs. More details are included in Section B of the Appendix.
@@ 4.2 Composed Video Retrieval Results
We provide a number of baselines for our new benchmark on WebVid-CoVR-Test. Table 2 summarizes these CoVR results. We first report the random chance performance in the first row. The rest of the table is split into two. The top block uses existing pretrained text and image encoders from BLIP (Li et al. 2022b) or CLIP (Radford et al. 2021) backbones without any finetuning. Models in the bottom block are finetuned on WebVid-CoVR. We report results with the composed query, as well as with the individual modalities. For combining modalities, we experiment with the simple average fusion baseline (Avg) when using frozen embeddings, and fusion with a randomly-initialized MLP or BLIP-pretrained cross-attention (CA) layers when finetuning. Note that the MLP fusion baseline is similar to Combiner (Baldrati et al. 2022) that concatenates the image and text embeddings from CLIP (or BLIP in CASE (Levy et al. 2024)), and is referred to as late fusion by CASE. For finetuning individual modalities, we train and test either with text-only query using the modification text, or with the visual-only image query. Finally, we experiment with using only the middle frame embedding or the weighted average of target video frame embeddings as explained in Section 3.3 (with the exception that visual-only experiments use equal weights due to not having access to the modification text for computing the scores).
Mode Pretraining CIRR FashionIQ
Method Data R@1 R@5 R@10 R@50 R@10 R@50
Train TIRG+LastConv (Vo ’19) $\dagger$ - 11.04 35.68 51.27 83.29 - -
TIRG (Vo ’19$\dagger$ - 14.61 48.37 64.08 90.03 - -
JVSM (Chen ‘20) - - - - - 11.90 26.60
TRACE w/BER (Jandial ‘20) - - - - - 22.57 46.19
VAL w/GloVe (Chen ‘20) - - - - - 24.15 46.61
CurlingNet (Yu ‘20$\ddagger$ - - - - - 25.90 51.01
RTIC-GCN (Shin ‘21$\ddagger$ - - - - - 28.18 53.09
CoSMo (Lee ‘21) - - - - - 26.58 52.31
DCNet (Kim ‘21) - - - - - 27.78 53.89
MAAF (Doods ‘20$\dagger$ - 10.31 33.03 48.30 80.06 24.30 48.80
MAAF-BERT (Doods ‘20$\dagger$ - 10.12 33.10 48.01 80.57 - -
MAAF-IT (Doods ‘20$\dagger$ - 09.90 32.86 48.83 80.27 - -
MAAF-RP (Doods ‘20$\dagger$ - 10.22 33.32 48.68 81.84 - -
CIRPLANT (Liu ‘21$\dagger$ - 19.55 52.55 68.39 92.38 18.87 41.53
SAC w/BERT (Jandial ‘22) - - - - - 29.08 54.70
FashionVLP (Goenka ‘22) - - - - - 34.27 62.51
ARTEMIS (Delmas ‘22) - 16.96 46.10 61.31 87.73 26.05 50.29
LF-CLIP (Baldrati ‘22) - - - - - 35.39 59.03
LF-BLIP (Levy 2023) - 20.89 48.07 61.16 83.71 25.75 43.98
Combiner (Baldrati ‘22) - 33.59 65.35 77.35 95.21 - -
CompoDiff (Gu ‘23) ST18M 22.35 54.36 73.41 91.77 40.98 62.23
CASE (Levy ‘23) - 48.00 79.11 87.25 97.57 - -
CASE (Levy ‘23) LaSCo 48.68 79.98 88.51 97.49 48.79 70.68
CASE (Levy ‘23) LaSCo+COCO 49.35 80.02 88.75 97.47 - -
CoVR-BLIP - 49.16 79.76 88.65 97.49 49.02 70.24
CoVR-BLIP (Ours) WebVid-CoVR 50.41 80.96 89.35 97.64 49.40 70.98

### v4

@@ 3.2 Our resulting WebVid-CoVR dataset
WebVid-CoVR: a large-scale CoVR training dataset. By applying the previously described pipeline to the WebVid2M dataset [12], we generate WebVid-CoVR, a dataset containing 1.6M CoVR triplets, which is significantly larger than prior datasets (see Table I). On average, a video lasts 16.8 seconds, a modification text contains 4.8 words, and one target video is associated with 12.7 triplets. We study the effect of the modification text length in Section D.5 of the Appendix. WebVid-CoVR is highly diverse with 131k distinct videos and 467k distinct modification texts. Examples of CoVR triplets from the WebVid-CoVR dataset are illustrated in Figure 3. These examples (along with additional ones included in Section E.3 ) demonstrate the diversity present in WebVid-CoVR, highlighting a wide range of content and variations in the modification texts. However, it is important to acknowledge that some noise naturally exists in the dataset, as shown in the bottom example of Figure 3, where the text does not describe the difference between the two videos due to both videos describing beautiful fields. We provide further analysis such as removal of inappropriate content, and dataset statistics of WebVid-CoVR in Section A of the Appendix.
WebVid-CoVR-Test: a new CoVR evaluation benchmark. Due to the noise in WebVid-CoVR, we manually annotate a small test set, dubbed WebVid-CoVR-Test, for evaluation. For this, we first repeat the data generation procedure described in Section 3.1, but on a different corpus of video-caption pairs. Specifically, we consider video-caption pairs from the WebVid10M corpus [12] that are not included in the WebVid2M dataset, resulting in a pool of 8 million video-caption pairs. This ensures that other models using WebVid2M for pretraining have not been exposed to any of the test examples. In the video pairs filtering stage, for each pair of captions, we here only keep one pair of videos (the one with the highest visual similarity). This results in 163k candidate triplets that could be used for testing purposes. We randomly sample 7k triplets that we use for validation and randomly sample 3.2k other triplets that we manually annotate as described below.
@@
Figure 4: Model architecture of CoVR-BLIP-2: The BLIP-2 [16] image encoder extracts visual features from the image query $q$. These visual features are combined with the text query $t$ (modification text) through the BLIP-2 image-grounded text encoder to obtain a multi-modal query embedding $f(q,t)$. To encode videos, $N$ frames are individually encoded with the BLIP-2 image encoder and its Q-Former, and aggregated via a weighted mean into a single video embedding $h(v)$. The goal of CoVR (video retrieval) is to maximize similarity between the multi-modal query $f(q,t)$ and the target video $h(v)$. During training, an additional caption retrieval loss $\mathcal{L_{c}}$ is defined between $f(q,t)$ and the target caption embedding $g(c)$. Note that for simplicity, we visualize one Q-Former block, but in practice there are 12 blocks as in [16]. To reduce the 32 tokens output by the Q-Former, we simply average them before computing cosine similarities. While each Q-Former is initialized from the BLIP-2 pretraining, they are finetuned on our CoVR/CoIR data. When training for CoIR, the target becomes a single image, removing the need for weighted mean. See Section 3.3 for more details.
@@ 3.3 Training on WebVid-CoVR
CoVR-BLIP-2 model architecture. Our model, illustrated in Figure 4, builds on a pretrained image-text model, BLIP-2 [16], an enhancement over the original BLIP [1]. The main architectural difference with the CoVR-BLIP approach in our preliminary work [58] is the use of BLIP-2 [16] instead of BLIP [1]. BLIP-2 introduces a lightweight Querying Transformer (Q-Former) with learnable queries for more efficient visual feature extraction. Like its predecessor, BLIP-2 is pretrained on a large dataset of image-caption pairs with three vision-language objectives: image-text contrastive learning, image-text matching, and image-conditioned language modeling. However, BLIP-2 is not trained for composed visual retrieval with both visual and text inputs. Therefore we adapt BLIP-2 to the CoIR/CoVR task as follows.
We use the BLIP-2 image encoder to encode the image query $q$ (which corresponds to the middle frame of the video in case of WebVid-CoVR). The resulting visual features and the modification text ($t$) are then forwarded to the BLIP-2 image-grounded text encoder together, which outputs a multi-modal embedding $f(q,t)\in\mathbb{R}^{d}$ where $d$ is the embedding dimension. Note that we compute mean across the 32 output tokens corresponding to the learnable query inputs of the Q-Former, after passing them through linear projection layer initialized from BLIP-2 text projection layer.
To retrieve a target video $v_{k}$ from a database of videos $V$, we compute embedding vectors for all gallery videos as follows. We uniformly sample $N$ frames from the video and compute a weighted mean of the BLIP-2 image embeddings to obtain the video embedding vector $h(v_{k})\in\mathbb{R}^{d}$. The weights $\{\omega\}^{N}$ are obtained by computing the similarity between the corresponding frame and the modification text using the pretrained BLIP-2 image and text encoders, respectively (introduced as ‘query scoring’ in [17] in the context of text-to-video retrieval). Using pretrained and frozen BLIP-2 embeddings allows us to precompute and store all these weights, which we refer to as $\{\omega\}^{N}=QS(\{v\}^{N},t)$ for query-scoring between each of the $N$ frames of the video $v$ and the text $t$.
@@
Training. In order to train on WebVid-CoVR, we use a contrastive learning approach [14, 15], as it has been shown to be effective to learn strong multi-modal representations from large-scale noisy data [2]. We make the following design choices. First, we create a training batch by sampling distinct target videos; and for each target video, we randomly sample an associated image-text query pair. Iterating over videos ensures that the same target video appears only once in a batch and maximizes the number of different target videos that can be used as negatives in contrastive learning. We show the benefit of this approach in Section 4.6 (Table X). Second, we employ HN-NCE [15] which increases the weight of most similar samples and uses as negatives all target videos $v_{j\in\mathcal{B}}$ in the batch $\mathcal{B}$. Formally, given a training batch $\mathcal{B}$ of triplets $(q_{i},t_{i},v_{i})$, we minimize the following loss:
$\mathcal{L}_{v}(\mathcal{B})=$ $-\sum_{i\in\mathcal{B}}\text{log}\left(\frac{e^{S_{i,i}/\tau}}{\alpha\cdot{e^{S_{i,i}/\tau}}+\sum_{j\neq i}{e^{S_{i,j}/\tau}w_{i,j}}}\right)$
$-\sum_{i\in\mathcal{B}}\text{log}\left(\frac{e^{S_{i,i}/\tau}}{\alpha\cdot{e^{S_{i,i}/\tau}}+\sum_{j\neq i}{e^{S_{j,i}/\tau}w_{j,i}}}\right)$ (1)
where $\alpha$ is set to 1, temperature $\tau$ is set to $0.07$, $S_{i,j}$ is the cosine similarity between the multi-modal embedding $f(q_{i},t_{i})$ and the target video embedding $h(v_{j})$, and $w_{i,j}$ is set as in [15] with $\beta=0.5$.
Composed caption retrieval as an additional loss term. In addition to using the video as a target, our approach also leverages the supervision from the caption corresponding to each video. This is possible in our training data because we in fact have 5 elements (video1, caption1, video2, caption2, modification text) for each data sample. This new loss term involves aligning the multi-modal query embedding $f(q,t)$ not only with the video embedding $h(v)$ but also with its descriptive caption $c$. To this end, we encode the caption into an embedding $g(c)$, using frozen BLIP-2 text encoding and define an additional contrastive loss term ($\mathcal{L}_{c}$). Therefore, our final loss can be expressed as:
$\mathcal{L}(\mathcal{B})=$ $\lambda_{v}\cdot\mathcal{L}_{v}(\mathcal{B})+\lambda_{c}\cdot\mathcal{L}_{c}(\mathcal{B}),$ (2)
where $\lambda_{v}$ and $\lambda_{c}$ are both set to 0.5. Here, $\mathcal{L}_{v}$ refers to the initial loss in Eq.(1) utilizing cosine similarity with the target video embedding $h(v_{j})$, and $\mathcal{L}_{c}$ employs the same loss structure but applies it to the cosine similarity with the target video caption embedding $g(c_{j})$.
@@ 4 Experiments
We first describe the experimental protocol including the datasets, evaluation metrics, and implementation details (Section 4.1). We then present the results of CoVR on our new video benchmark (Section 4.2). Additionally, we introduce CC-CoIR, a new large-scale CoIR training dataset derived with a similar methodology, from the Conceptual Captions (CC) dataset (Section 4.3). We show transfer results of CoIR on standard image benchmarks, together with an extensive state-of-the-art comparison (Section 4.4). We further provide a comparison by training on other automatic triplet datasets (Section 4.5). Finally, we provide ablations on our key components such as the caption retrieval loss and data scale. (Section 4.6) and we illustrate qualitative examples in Section 4.7).
@@ 4.1 Experimental setup
Datasets. WebVid-CoVR is our proposed training CoVR dataset, and WebVid-CoVR-Test is our new CoVR benchmark, both presented in Section 3.2. For pre-training on CoIR datasets, we use WV-CC-CoVIR, a combination of both WebVid-CoVR and CC-CoIR which is a new large-scale CoIR training dataset derived from the Conceptual Captions dataset, that we generate using the same methodology as WebVid-CoVR. See Section 4.3 for more details.
CIRR [8] is a manually annotated CoIR dataset that contains open-domain natural images from NLVR2 [59], comprising 36.5k queries annotated on 19k images. CIRR includes two evaluation protocols: a standard one with the entire validation set as the search gallery, and a fine-grained subset, where the search space is a subgroup of six images similar to the query image (based on pretrained ResNet15 feature distance). The dataset is divided into training, validation, and testing splits with 28225/16742, 4181/2265 and 4148/2178 queries/images, respectively.
FashionIQ [9] is another CoIR dataset that contains images of fashion products, divided into three categories of Shirts, Dresses, and Tops/Tees. The query and target images were automatically paired based on title similarities (crawled from the web), and modification texts were then manually annotated. This dataset consists of 30k queries annotated on 40.5k different images. It is divided into training and validation splits with 18000/45429 and 6016/15415 queries/images, respectively.
CIRCO [4] is an open-domain dataset for CoIR, specifically designed for zero-shot CoIR tasks, as there is no specific training split provided for this dataset. It is unique in its inclusion of multiple ground truths for each query, with an average of 4.53 ground-truth images per query. This feature allows for a more reliable and robust evaluation using the mean Average Precision (mAP) metric. CIRCO is divided into a validation split (220 queries) and a test split (800 queries). The evaluation protocol uses all 120,000 images from the COCO dataset as its gallery set.
Evaluation metrics. Following standard evaluation protocols [8], we report the video retrieval recall at rank 1, 5, 10, and 50. Recall at rank k (R@k) quantifies the number of times the correct video is among the top k results. MeanR denotes the average of R@1, R@5, R@10, and R@50. Higher recall means better performance.
Implementation details and environmental costs. For our MTG-LLM, we use LLaMA 7B model [10] that we finetune for one epoch with an initial learning rate of $3\mathrm{e}{-5}$. For our CoVR model, we use the BLIP-2 with ViT-G/14 [60] at 364 pixels finetuned for text-image retrieval on COCO and freeze the ViT for computational efficiency. We train our CoVR model on WebVid-CoVR for 5 epochs with a batch size of 2048 and an initial learning rate of $1\mathrm{e}{-5}$. To finetune on CIRR/FashionIQ, we train for 6 epochs with a batch size of 2048/1024 and an initial learning rate of $1\mathrm{e}{-4}$. We set hyperparameters based on the validation curve of WebVid-CoVR. Experiments are conducted on 4 NVIDIA A100-SXM4-80GB GPUs. The experiments conducted in this study incurred an environmental cost of approximately 49kg of $CO_{2}$ emissions. More details are included in Section C of the Appendix.
TABLE II: Benchmarking on the WebVid-CoVR-Test set: We observe that using both the visual and text input modalities performs better than individual modalities alone, both with/without finetuning on WebVid-CoVR (shown at the top/bottom of the table, respectively). When using pretraining models without finetuning, we apply average fusion (Avg) for the embeddings. BLIP performs slightly better than CLIP on this benchmark. Finetuning on WebVid-CoVR brings significant benefits. In this case, fusing with the pretrained cross-attention (CA) from BLIP is more effective than training a randomly-initialized MLP fusion as done in [5]. Moreover, using multiple frames to embed the target video brings further improvements over using the middle frame. The first row represents the random performance.
@@
modalities Fusion Backbone R@1 R@5 R@10 R@50
Random - - - 00.08 00.23 00.35 01.76
Text - BLIP-2 18.74 38.11 47.97 67.21
Not finetuned on Visual - BLIP-2 33.10 59.55 69.80 88.85
WebVid-CoVR Visual + Text Avg CLIP 44.37 69.13 77.62 93.00
Visual + Text Avg BLIP 45.46 70.46 79.54 93.27
Visual + Text Avg BLIP-2 45.66 71.71 81.30 94.80
Text - BLIP-2 23.32 46.21 56.22 78.36
Visual - BLIP-2 36.03 64.24 74.77 92.64
Visual + Text MLP CLIP 50.86 77.46 85.32 96.75
Finetuned on Visual + Text MLP BLIP 50.59 74.65 83.57 95.46
WebVid-CoVR Visual + Text MLP BLIP-2 51.88 79.38 86.46 97.42
Visual + Text CA BLIP 55.95 81.22 89.05 98.08
Visual + Text CA BLIP-2 59.82 83.84 91.28 98.24
@@ 4.2 Composed video retrieval results
We provide a number of baselines for our new benchmark on WebVid-CoVR-Test. Table II summarizes these CoVR results. We first report the random chance performance in the first row. The rest of the table is split into two. The top block uses existing pretrained text and image encoders from BLIP [1], BLIP-2 [16] or CLIP [2] backbones without any finetuning. Models in the bottom block are finetuned on WebVid-CoVR. We report results with the composed query, as well as with the individual modalities. For combining modalities, we experiment with the simple average fusion baseline (Avg) when using frozen embeddings, and fusion with a randomly-initialized MLP or BLIP-pretrained cross-attention (CA) layers when finetuning. Note that the MLP fusion baseline is similar to Combiner [5] that concatenates the image and text embeddings from CLIP (or BLIP in CASE [7]), and is referred to as late fusion by CASE. For finetuning individual modalities, we train and test either with text-only query using the modification text, or with the visual-only image query. Finally, we experiment with using the weighted average of target video frame embeddings as explained in Section 3.3 (with the exception that visual-only experiments use equal weights due to not having access to the modification text for computing the scores).
We make several conclusions. (i) Combining both visual and text modalities yields better performance than the models with individual modalities. This result highlights that our new CoVR benchmark requires paying attention to both modalities. (ii) Visual-only outperforms text-only suggesting that the video pairs automatically mined through their caption similarity indeed exhibits visual similarity, and that the image captures the target video better than the modification text. (iii) Finetuning on WebVid-CoVR obtains substantial improvements over using pretrained and frozen embeddings. (iv) When finetuning, fusion with BLIP-2 cross-attention (CA) performs better than the MLP fusion. (v) Results with the BLIP-2 backbone are higher than those with CLIP or BLIP. We analyze the effect of the number of frames in Section D.6.
4.3 Effect of training with CC-CoIR
We apply the same automatic triplet generation procedure to the Conceptual Captions dataset [13] (CC3M) resulting in the formation of CC-CoIR. This new dataset comprises 3.3M CoIR triplets, utilizing the diverse and contextually rich content of CC3M [13]. In contrast to WebVid-CoVR, which is video-based, CC-CoIR exclusively incorporates images, adding a different modality to our training material. The average length of modification texts in this dataset is 24.65, featuring 130k unique images and 28k distinct modification texts.
Table III presents the results when training with our newly introduced CC-CoIR dataset. When used independently for pretraining (zero-shot) we observe the following: (a) CC-CoIR outperforms WebVid-CoVR on the CIRR dataset, (b) CC-CoIR has similar performance than WebVid-CoVR in FashionIQ, (c) CC-CoIR has lower performance for the CIRCO benchamark, and (d) CC-CoIR performance has good zero-shot results on our video retrieval test set WebVid-CoVR-Test. Note that WebVid-CoVR-Test constitutes a zero-shot setting only when trained using the CC-CoIR pretraining, whereas pretraining on WebVid-CoVR means the model has seen samples from the same data distribution during training. By combining both WebVid-CoVR and CC-CoIR into a unified pretraining dataset WV-CC-CoVIR, we observe a slight improvement on the image-based datasets while experiencing a minor drop on WebVid-CoVR-Test compared to using only WebVid-CoVR. We hypothesize that this minor drop is due to a domain gap with CC-CoIR, which is included in WV-CC-CoVIR. We therefore opt to use the jointly pretrained model going forward, as it leverages the strengths of both data sources to perform well across different datasets (i.e., best average recall across all datasets).
TABLE III: Training with CC-CoIR: We compare the performance metrics when pretraining with various combinations of WebVid-CoVR and CC-CoIR datasets. We observe that in the zero-shot setting, combining both datasets yields the best overall results.
WebVid- CC- WebVid-CoVR-T. CIRR FashionIQ CIRCO
CoVR CoIR R@1 R@1 R@10 mAP@5
53.83 43.35 36.49 23.50
59.82 41.42 36.81 28.88
57.71 43.74 38.15 28.29
4.4 State-of-the-art comparison on CoIR benchmarks
While our focus is video retrieval, we also experiment with transferring our CoVR models to image retrieval tasks on standard CoIR benchmarks. We define zero-shot CoIR as not using any manually annotated CoIR triplet for training. To perform zero-shot CoIR, we directly apply our model, which has been trained on our automatically generated WV-CC-CoVIR dataset, to CoIR tasks. In addition to zero-shot CoIR, for the datasets that have training split (CIRR and FashionIQ), we investigate how our model performs when finetuned on the training set of the downstream benchmark.
TABLE IV: State-of-the-art comparison on the CIRR and FashionIQ test sets: Our model benefits from training on our datasets in the zero-shot setting, and in the finetuning setting where it performs competitively. $\dagger$ denotes results reported by [8]. For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. CC3M denotes Conceptual Captions 3M [13], ST18M denotes  [6], and COCO denotes [61].
Mode CIRR FashionIQ
Pretraining Recall@K $\text{R}_{\text{subset}}$@K Dress Shirt Toptee Average
Method Year Data K=1 K=5 K=10 K=50 K=1 K=2 K=3 R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
Train TIRG [3$\dagger$ 2019 - 14.61 48.37 64.08 90.03 22.67 44.97 65.14 14.87 34.66 18.26 37.89 19.08 39.62 17.40 37.39
TIRG+LastConv [3] $\dagger$ 2019 - 11.04 35.68 51.27 83.29 23.82 45.65 64.55 - - - - - - - -
CurlingNet [62$\ddagger$ 2019 - - - - - - - - 26.15 53.24 21.45 44.56 30.12 55.23 25.90 51.01
JVSM [63] 2020 - - - - - - - - 10.70 25.90 12.00 27.10 13.00 26.90 11.90 26.60
TRACE w/BER [64] 2020 - - - - - - - - 22.70 44.91 20.80 40.80 24.22 49.80 22.57 46.19
VAL w/GloVe [65] 2020 - - - - - - - - 22.53 44.00 22.38 44.15 27.53 51.68 24.15 46.61
MAAF [66$\dagger$ 2020 - 10.31 33.03 48.30 80.06 21.05 41.81 61.60 23.80 48.60 21.30 44.20 27.90 53.60 24.30 48.80
MAAF-BERT [66$\dagger$ 2020 - 10.12 33.10 48.01 80.57 22.04 42.41 62.14 - - - - - - - -
MAAF-IT [66$\dagger$ 2020 - 09.90 32.86 48.83 80.27 21.17 42.04 60.91 - - - - - - - -
MAAF-RP [66$\dagger$ 2020 - 10.22 33.32 48.68 81.84 21.41 42.17 61.60 - - - - - - - -
RTIC-GCN [67$\ddagger$ 2021 - - - - - - - - 29.15 54.04 23.79 47.25 31.61 57.98 28.18 53.09
CoSMo [68] 2021 - - - - - - - - 25.64 50.30 24.90 49.18 29.21 57.46 26.58 52.31
DCNet [20] 2021 - - - - - - - - 28.95 56.07 23.95 47.30 30.44 58.29 27.78 53.89
CIRPLANT [8$\dagger$ 2021 - 19.55 52.55 68.39 92.38 39.20 63.03 79.49 17.45 40.41 17.53 38.81 61.64 45.38 18.87 41.53
SAC w/BERT [69] 2022 - - - - - - - - 26.52 51.01 28.02 51.86 32.70 61.23 29.08 54.70
FashionVLP [70] 2022 - - - - - - - - 32.42 60.29 31.89 58.44 38.51 68.79 34.27 62.51
ARTEMIS [19] 2022 - 16.96 46.10 61.31 87.73 39.99 62.20 75.67 27.16 52.40 21.78 43.64 29.20 53.83 26.05 50.29
LF-CLIP [5] $\dagger$ 2022 - 33.59 65.35 77.35 95.21 62.39 81.81 92.02 31.63 56.67 36.36 58.00 38.19 62.42 35.39 59.03
LF-BLIP [5, 7] $\dagger$ 2022 - 20.89 48.07 61.16 83.71 50.22 73.16 86.82 25.31 44.05 25.39 43.57 26.54 44.48 25.75 43.98
Combiner [5] 2022 - 33.59 65.35 77.35 95.21 62.39 81.81 92.02 31.63 56.67 36.36 58.00 38.19 62.42 35.39 59.03
CompoDiff [6] 2023 ST18M+LAION2B 32.39 57.61 77.25 94.61 67.88 85.29 94.07 38.39 51.03 41.68 56.02 45.70 57.32 39.81 51.90
CASE [7] 2024 LaSCo [7] 48.68 79.98 88.51 97.49 76.39 90.12 95.86 47.44 69.36 48.48 70.23 50.18 72.24 48.79 70.68
CASE [7] 2024 LaSCo+COCO 49.35 80.02 88.75 97.47 76.48 90.37 95.71 - - - - - - - -
CoVR-BLIP [58] 2024 WebVid-CoVR 49.69 78.60 86.77 94.31 75.01 88.12 93.16 44.55 69.03 48.43 67.42 52.60 74.31 48.53 70.25
CoVR-BLIP-2 2024 - 50.87 80.80 88.84 98.00 76.70 90.31 95.45 45.25 68.86 49.95 69.95 51.37 72.56 48.86 70.46
CoVR-BLIP-2 2024 CC-CoIR 50.63 81.04 89.35 98.15 76.53 90.43 96.00 46.49 69.31 51.67 70.49 51.53 73.07 49.90 70.96
CoVR-BLIP-2 2024 WV-CC-CoVIR 50.43 81.08 88.89 98.05 76.75 90.34 95.78 46.53 69.60 51.23 70.64 52.14 73.27 49.96 71.17